40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed Oct 30 01:56:38 2013, tara, DailyProgress, optic, table work oldPBS.jpgnewPBS.jpg
    Reply  Thu Oct 31 00:02:17 2013, tara, DailyProgress, Electronics Equipment, TTFSS offset.pngD040105C.png
       Reply  Thu Oct 31 18:10:07 2013, Evan, DailyProgress, Electronics Equipment, TTFSS 
       Reply  Sun Nov 3 19:37:32 2013, rana, DailyProgress, Electronics Equipment, TTFSS 
          Reply  Tue Nov 5 01:11:15 2013, tara, DailyProgress, Electronics Equipment, TTFSS 
             Reply  Wed Nov 6 01:14:58 2013, tara, DailyProgress, Electronics Equipment, TTFSS beat_2013_11_05.pngbeat_2013_11_05.fig
                Reply  Thu Nov 7 05:08:13 2013, tara, DailyProgress, NoiseBudget, photothermal noise in SiO2/Ta2O5 beat_2013_11_07.pngbeat_2013_11_07.fig
                   Reply  Wed Nov 20 18:19:01 2013, tara, DailyProgress, NoiseBudget, photothermal noise in SiO2/Ta2O5 beat_compare.pngbeat_compare.fig
Message ID: 1377     Entry time: Thu Oct 31 00:02:17 2013     In reply to: 1376     Reply to this: 1378   1380
Author: tara 
Type: DailyProgress 
Category: Electronics Equipment 
Subject: TTFSS 

Evan found that when common gain is changed, DC offset also changes as well. I'm still looking into the problem.

 

D040105C.png

a part of schematic, the driving signal was sent in through test port (the switch was flipped from off to test), so the signal came through PD line in this page.

 We still cannot lock RCAV with TTFSS, so I'm checking the box 2009007 (#7).

  • The modifications I did two days ago were 1) adding a push switch for gain reduction and 2) replacing one resistor. The rests were changed before we got the TTFSS.
  • I checked the TF between TP1 and TP5, it works as it should be (20log( 390/100) ~ 12 dB). So this stage does not have any problem. I checked both TF and time domain signals. So the modifications are ok.
  •  Note, when I measured the TF of TP1/TP4 or TP5/TP4, the signals oscillated and became very noisy. I don't understand why, but this problem disappeared when I used TP4 and out2. Both boxes (#5,#7) have this same problem.

Common Gain - DC offset problem

  • When Common gain is increased (CG signal to U2A chip), there is an offset observed in TP4 which is after the variable gain stage. Both boxes behave similarly. <- This surprises me, as we haven't seen this (or haven't noticed this) before.
  • I checked if the offset varied with the input drive or not. I changed the input from 20mV to 40mV, with constant gain = 1000(25dB). The behavior is nonlinear (see the plot below). I checked this only on box#5.

offset.png

DC offset vs input drive. DC offset is calculated from (Vmax + Vmin) /2  from a sinusoidal signal input. The signal was taken from TP4. The behavior is very non linear and it is impossible to make a table for an appropriate offset level vs common gain setting.

 What to do next?

  • This seems to be an important clue about why the loop behaves badly when common gain is increased. From today test, both boxes behave the same, so I think it might be the chips' problem.
  • Fortunately, we can use the offset adjustment(OS)
  • to cancel the offset introduced by the common gain, but we might need to add a port some where (we might be able use fast mon channels during laser scan).  So that when we increase the gain, we can adjust the offset accordingly.
  • Box#7 that I modified should be working. I don't know why I could not lock the laser before, more checking has to be done.

 

ELOG V3.1.3-